Each of ten tickets is marked with a different number from 1 to 10 and put in a box. IF you draw a ticket at random from the box, what is the probability that you will draw 4, 8, or 6?
Answer by Fombitz(13823)

P(4)=1/10
P(6)=1/10
P(8)=1/10
The events are mutually exclusive so,
P(4 or 6 or 8)=P(4)+P(6)+P(8)=1/10+1/10+1/10=3/10.